OpenStudy (dakid88):

A grocer sells oranges in bags. Each empty bag weighs 2.5 ounces and each orange weighs 7.2 ounces. Khalil buys a bag of oranges that weighs 60.1 ounces. Which equation can be used to find x, the number of oranges in the bag?

OpenStudy (dakid88):

A.\[7.2+x=60.1\] B.\[2.5x+7.2=60.1\] C.\[60.1-x=2.5+7.2\] D.\[2.5+7.2x=60.1\]

OpenStudy (dakid88):

@random231 @Coolguy34 @CoolDude12 @Adi3

random231 (random231):

weight of each orange X no of oranges bought + weight of bag = total weight

OpenStudy (dakid88):

thxs